will a tutor please help me with this problem,
Ann's present age exceeds Bob's age by 25 years. In 15 years Ann will be twice as old as bob.Find both ages
Ann now-
Bob now-
Ann in 15-
Bob in 15-

Answer by sdmmadam@yahoo.com(530)

Ann's present age exceeds Bob's age by 25 years. In 15 years Ann will be twice as old as bob.Find both ages
Ann now-
Bob now-
Ann in 15-
Bob in 15-
Let Ann be A yrs old now
and let Bob be B yrs old now
Ann's present age exceeds Bob's age by 25 years
That is A = B+25 ----(1)
In 15 years Ann will be twice as old as bob.
In 15 years
Ann will be (A+15) yrs old
and Bob will be (B+15) yrs old
And then Ann will be twice as old as bob.
Therefore (A+15)= 2(B+15)
A+15 = 2B +30
A-2B = 30-15
A-2B = 15 ----(2)
Using (1), that is A = B+25
(B+25)-2B = 15
(B-2B) = 15-25
-B = -10
Therefore B = 10
Putting B = 10 in (1)
A = B+25 = 10+25 = 35
Answer: Ann's present age is 35 yrs
and Bob's present age is 10 yrs
In 15 yrs
Ann will be (35+15) = 50 yrs
and Bob will be (10+15) = 25 yrs
Verification: In 15 years Ann should be twice as old as bob
which is of course true as 50 is twice 25.